Two point charges q1=4uc& q2=9uc are placed 20cm apart.the electric field due to them will be zero on the line joining them at a distance of
1
Expert's answer
2020-02-17T09:18:13-0500

For the equilibrium:


"\frac{q_1}{(x)^2}=\frac{q_2}{(d-x)^2}"

"\frac{4}{(x)^2}=\frac{9}{(20-x)^2}"

"x=8\ cm"

The electric field due to them will be zero on the line joining them at a distance of 8 cm from the first charge.